### v2

Abstract
Service robots work in a changing environment habited by exogenous agents like humans. In the service robotics domain, lots of uncertainties result from exogenous actions and inaccurate localisation of objects and the robot itself. This makes the robot task scheduling problem challenging. In this article, we propose a benchmarking framework for systematically assessing the performance of algorithms scheduling robot tasks. The robot environment incorporates a map of the room, furniture, transportable objects, and moving humans. The framework defines interfaces for the algorithms, tasks to be executed, and evaluation methods. The system consists of several tools, easing testing scenario generation for training AI-based scheduling algorithms and statistical testing. For benchmarking purposes, a set of scenarios is chosen, and the performance of several scheduling algorithms is assessed. The system source is published to serve the community for tuning and comparable assessment of robot task scheduling algorithms for service robots. The framework is validated by assessment of scheduling algorithms for the mobile robot executing patrol, human fall assistance and simplified pick and place tasks.
keywords:
[…]
{graphicalabstract}[Uncaptioned image]
{highlights}
Goal: Comparable evaluation of various algorithms scheduling diverse robot tasks.
Method: Kinematic simulation executed in configured or generated scenarios and environments.
Result: The framework teaches an RL-based algorithm, and other algorithms are compared.
Conclusion: The comparison shows algorithm performance for specific scenarios and tasks.
1 Introduction
1.1 Problem statement
Humans have honed the ability to manage and prioritise multiple tasks for centuries. Today, the challenge lies in transferring this capability to robots. This task has become increasingly important due to the growing need for robotisation in various domains facing staff shortages. Thus, robots ability to multitask efficiently becomes critical, and the proposed scheduling algorithms should be assessed systematically.
Solving the problem of robot task management requires adapting task scheduling solutions to real-world environments characterised by dynamic changes [8], uncertainty [13], and interactions with exogenous actors such as humans [12, 18]. This presents a significant challenge, as changing environments affect the schedules and action plans required to perform the requested tasks. For safety and effectiveness reasons, task scheduling algorithms should be rigorously tested in simulated environments [3, 16]. Evaluations of these algorithms may involve statistical analysis or proving their optimality. The latter is especially difficult in changing, inhabited environments, particularly for neural-network-based scheduling algorithms [15].
Task scheduling for mobile robots in dynamic environments has been extensively studied, with researchers exploring both traditional and AI-based methods to address the inherent challenges. Classical approaches often rely on heuristic or optimization-based strategies. However, these methods struggle with scalability and adaptability in environments with frequent changes. Modern AI-driven methods, such as reinforcement learning, have shown promise in addressing dynamic scheduling. For example, [30] review studies focused on Reinforcement Learning techniques that have been used for dynamic task scheduling. [33] argue that a Hybrid Q-learning approach allows for maintaining optimal scheduling. Furthermore, they state their reinforcement learning method is resilient to non-deterministic environments. They validate the algorithm in a robotic arm environment, and in the future, they want to focus on motion planning, collision avoidance, and multi-agent algorithms. Similarly, [36] proposed a deep multi-agent reinforcement learning approach to optimize task scheduling in human-robot collaboration, showcasing its effectiveness in dynamic environments. Yet, these techniques often require extensive training data and may falter in highly unpredictable scenarios. Simulation tools like Gazebo and frameworks such as ROSPlan [5] have facilitated the testing of scheduling algorithms, but they are limited in their ability to capture the complexities of human-robot interactions and real-world uncertainties. Despite these advancements, benchmarking service robot scheduling algorithms in such environments remains an open problem, as existing tools and methods fall short in addressing the unpredictable and stochastic nature of real-world settings, and transparent comparison of the methods is difficult. Thus, a framework for comparable benchmarking of the algorithms is mandatory.
Although current frameworks and benchmarking systems provide essential tools for evaluating task scheduling algorithms, they often fall short in addressing the complexities of dynamic and uncertain environments in the field of service robotics. The framework introduced in this work aims to bridge this gap by offering a structured approach to training and benchmarking scheduling algorithms. The proposed framework explicitly accounts for real-world uncertainties and facilitates fair performance comparisons across diverse scenarios. The framework is adaptable, supports various environments and algorithms, and allows for systematic and repeatable testing against configurable scenarios. Importantly, the framework ensures that performance scores are comparable between different algorithms and scenarios. In addition, it provides tools for generating test cases and scenarios, which can be used to train AI-based scheduling algorithms. Additionally, this article introduces a SysML-based [35] domain-specific language (DSL) designed to support the presentation, documentation, and analysis of benchmarking systems developed with the proposed system. The DSL follows the Model-based Systems Engineering (MBSE) method used broadly for systems analysis, design verification and validation in the V-model [6] workflow. The framework’s capabilities are illustrated through the use case diagram Fig. 1. As target users of the framework, we recognise the scientists developing scheduling algorithms and mobile manipulator integrators. The algorithm developers can use the framework to test their solutions against other algorithms and previous versions of their work in a comparable environment. The mobile manipulator integrators can mimic a use-case-specific environment in the framework and validate which scheduling algorithms best suit the tasks and environment of the robot application.
Figure 1: Use cases of the proposed benchmarking system
The framework should ease analysing how the scenario’s change influences the algorithm’s score. This article investigates how such a system should be structured, how it works, and how to use it to receive comparable results of different scheduling algorithms evaluated in various scenarios.
The best-assessed scheduling algorithm must be integrated into the end product— robot control system. For the architecture-level integration, we have proposed the Simulation-Physical Modeling Language [9], and for the implementation-level integration, we have published the ROS [^1] Robot Operating System [19, 34]-based TaskER framework [10] that models the structure and behaviour of a multitasking robot. TaskER manages safe interruptions between different task contexts, allowing the robot to conduct safety-critical actions of the interrupted task before executing the switch to the next task.
The contribution of this article can be summarized as follows:
  • a comprehensive framework for training and benchmarking robot task scheduling algorithms, explicitly addressing real-world uncertainties in dynamic and human-inhabited environments. It supports systematic, repeatable testing across configurable scenarios while ensuring comparable performance evaluation.
  • a dedicated domain-specific language for presenting, documenting, and analyzing benchmarking systems, enhancing transparency and standardization in the development and assessment of scheduling algorithms.
  • an Open-Source and Practical Validation: The framework, validated through scenarios involving patrol, human fall assistance, and pick-and-place tasks, is published as an open-source tool, promoting community collaboration and the development of AI-based scheduling algorithms.
In the next section 2, we summarise the state-of-the-art solutions to the problem of benchmarking task scheduling algorithms. In the following sections, we introduce the architecture of the proposed benchmarking system 3.1, the scope of possible benchmarking scenarios 3.2, the configuration of the system for testing specific algorithms and robots 3.4, and finally, we describe the developed tools for generating test-cases 3.5. In section 4, we describe the configuration of the example system that evaluates basic and neural-network-based algorithms. The neural network algorithms were taught on the test cases generated by the proposed tools. Finally, the evaluation results are presented and discussed in 4.2.
2 Related work
2.1 Simulators for developing and benchmarking in robotics
Simulators play a crucial role in developing and benchmarking algorithms in the robotics domain. The Stage simulator, as discussed by [14], provides a simplified platform for benchmarking navigation systems in static environments. However, its project has been inactive for a long time, making its utility limited by the challenges of integrating recent algorithms. Additionally, the review of physics-based simulators by [23] emphasizes the importance of simulation accuracy in testing robotic controllers and the potential benefits of simplifying simulations for high-level task scheduling problems. The varying performance of simulators across different robotic domains, as noted by [7], further underscores the need for tailored benchmarking systems that account for the specific physical effects relevant to each application.
Simulation is a crucial development tool for robotics, which allows avoiding challenges of learning on the real robot, especially when learning-based techniques are applied [17]. This approach avoids the time-consuming experiments on real platforms and mitigates the risk of damaging the real robots. Moreover, the utilization of parallel architecture of the Graphical Processing Units enables simulating many robots at the same time and significantly reduces the training time [28]. The new PhysX-based Orbit platform standardizes the benchmarking manipulation and robot locomotion tasks [25].
2.2 Benchmarking task scheduling
The problem of robot task scheduling in dynamic and uncertain environments has garnered significant attention in the field of robotics. In the general domain of task scheduling, several frameworks and benchmarking systems have been proposed. One notable framework is introduced by [27], which leverages large language models (LLMs) for task division and tool selection. The framework evaluates the performance of three LLMs by measuring the success rate and the order of tool usage in completing subtasks. This approach demonstrates the potential of LLMs in enhancing task scheduling through intelligent tool selection, although its primary focus is on utilising language models rather than the broader context of robot task scheduling. LLMs are also used to interpret vocal requests of tasks and identify their parameters [32]. Another significant contribution is TASKOGRAPHY [1], a large-scale benchmark designed to evaluate robotic task planning over 3D scene graphs (3DSGs). This work also introduces SCRUB, a planner-agnostic strategy for adapting 3DSGs to improve planning performance, and SEEK, a procedure that enhances learning-based planners’ ability to exploit 3DSGs. These tools are particularly relevant for environments requiring detailed scene descriptions, although they focus more on planning than scheduling tasks in a dynamic environment.
The survey by [4] provides a comprehensive overview of various task scheduling metrics, offering a foundational understanding of the criteria used to assess scheduling algorithms. This work, along with the review of metrics and benchmarking for parallel job scheduling by [11], highlights the diversity of metrics applicable to scheduling, though these works primarily concentrate on static environments. In the domain of task and motion planning (TAMP), [21] proposes a platform-independent evaluation method, which is crucial for assessing the generalizability of scheduling algorithms across different robotic platforms. Similarly, the work by [27] offers a structured framework for LLM-based AI agents, focusing on the execution of inference processes, but again, with a specific focus on language models rather than broader task scheduling in robotics.
For benchmarking dynamic job scheduling, [24] presents the Dynamic Job Scheduler Benchmark (DJSB), a tool for comparing performance metrics across different scenarios. This tool, which allows for testing various resource management strategies, is particularly relevant for applications where resource availability and task demands change over time. The framework proposed by [2] addresses the PET-aware task-to-core scheduling problem, offering evaluation procedures and benchmarks for comparing scheduling algorithms. This work, along with the graph-based scheduling algorithm benchmarking by [20], contributes valuable methodologies for assessing the efficiency and performance of scheduling algorithms, particularly regarding resource utilization and execution time.
In summary, while existing frameworks and benchmarking systems offer valuable tools and methodologies for evaluating task scheduling algorithms, there remains a need for a comprehensive system that addresses the unique challenges posed by dynamic and uncertain environments in the service robotics domain. Our proposed framework seeks to fill this gap by providing a systematic approach to training and benchmarking scheduling algorithms, incorporating real-world uncertainties and enabling fair comparisons across different scenarios. Besides the above core functional contribution, this article provides a domain-specification language (DSL) for the presentation, documentation, and analysis of the benchmarking systems that are built using the proposed framework. The DSL is expressed in Systems Modeling Language (SysML) [35].
3 The framework for benchmarking and training scheduling algorithms
3.1 The architecture
TaBSA— the framework for Training and Benchmarking Scheduling Algorithms is a metamodel that, by configuration, gives a specific benchmarking system named TaBSA System. The TaBSA System is organised in a way presented in the SysML block definition diagram in Fig. 2. The diagram is the part of the metamodel for the variety of particular TaBSA System configurations (realisations). For each block, a dedicated stereotype was introduced to simplify the specification of these realisations. The TaBSA System addresses the use cases [UCX] from Fig. 1, by, a.o., reflecting the entities from these use cases as dedicated stereotypes: <<TaBSASystem>> – [UC1] and [UC2], <<Scenario>> – [UC1.2], <<Robot>> – [UC1.2.1], <<Task>> – [UC1.2.2], <<Environment>> – [UC1.2.3], <<DecAgent>> – [UC1.3], <<EvalFunction>> – [UC1.4].
Figure 2: TaBSA system structure
All system elements that can constitute the variety of configurations are composed (filled rhombus) into <<TaBSASystem>>. The set of the elements for the current configuration (it addresses configurability) of the system is represented by the aggregations (empty rhombus) into the corresponding blocks.
The <<TaBSASystem>> general behaviour is also a part of the metamodel. It is depicted in activity diagram 3. It provides a scheme for repeatable execution of configurable scenarios [UC1]. Activity [ACT1] addresses configurability, [ACT2] in particular addresses teaching and a step of the task execution [UC2], [ACT3] evaluates the latest task change decision [UC1.4]. A detailed description of the particular operations mentioned in the diagram is presented in the following part of the article.
Figure 3: TaBSA system main operation
3.2 Scope of benchmark scenarios
Apart from using a specific <<Robot>>’s model, the <<Scenario>> defines important aspects necessary for testing and training task management agents. These are the ’tasks’ list and the ’environment’. The <<Scenario>> also defines when the simulation starts and ends. Its capabilities can be extended using prepared plugins. The scenario’s most important operations are presented in the form of pseudocode in Fig. 4.
Figure 4: Pseudocodes for some operations of Scenario, Task, and EvalFunction
The ’tasks’ list contains every <<Task>> the robot should work on during the specific training session. We can define the different types of <<Task>>s our robot is supposed to perform as long as they can be represented in our restricted environment, as per [UC1.2.2]. This includes, a.o., simple operations such as navigation, picking, and placing of objects. Once we decide, we can fill the list with the selected amount for each type of <<Task>>. Each <<Task>> must possess the ’request_time’ property that represents at what time it is presented to the robot. The <<Task>>’s operations include is_called() and is_completed(). Apart from ’tasks’, the scenario also possesses the ’jobs’ list. This list contains solely the <<Task>>s that have already been called (is_called() operation returns True), so it can never contain a <<Task>> that can’t be found on the ’tasks’ list.
The <<Scenario>> defines the ’environment’ (of <<Environment>> stereotype) consisting of a flat map with walls, door openings, and furniture (simple shapes blocking the movement), items placed on the furniture, and moving humans represented by footprints of their legs, as per use case [UC1.2.3]. Simulated humans move on the map using simple kinematic equations and PRM algorithm for path planning, and their behaviour defines what happens after they reach the end of their path. They can circle between two points, choose a new destination, instantly change their position and choose a new destination, or just disappear. Humans are dynamically embedded on the map to make it possible for the robot to consider them during navigation. This represents the system incorporating external sensors for human detection. Manipulation targets are described solely by objects’ location. These provide a highly simplified representation of a real robot’s environment but are sufficient to evaluate <<Task>> execution order and allow training of neural network-based algorithm in a reasonable time. Due to its simplicity, it also satisfies the use case [UC2.2].
Table 1: Task operations as called by different components of our example benchmarking system — Mobile Manipulator System (for example pseudocodes see Fig. 4)
Operation Called by
is_called() <<Scenario>> - to determine when to add the <<Task>> to ’jobs’ list
is_completed() <<Scenario>> - to determine when to remove <<Task>> from ’jobs’ list
DQN Eval (Fig. 7) - to determine when the <<Task>> is completed and proper reward should be granted
Statistic Eval (Fig. 7) - to determine when the <<Task>> is completed to calculate time difference between
     planned and final completion
is_alive() DQN Eval (Fig. 7) - to determine if the <<Task>> died and therefore <<Scenario>> should terminate
Statistic Eval (Fig. 7) - to determine if the <<Task>> died and therefore <<Scenario>> should terminate
work_for() Mobile Manipulator (Fig. 7) - inside it’s execute_step() operation, it is passed as argument by the
     <<Scenario>>’s work() operation
estimate_duration() <<Scenario>> - to update <<Task>>’s estimated durationbased on it’s own properties. It is called
     by update_job() operation
3.3 Task features
Among the operations of <<Task>>, the most important is work_for(), whose implementation determines how the <<Task>> is executed. Some <<Task>>s may be critical, and exceeding their deadline may cause serious consequences. Thus, we define deathtime attribute of <<Task>>. Deathtime equals:
$deathtime=deadline+max\_delay.$ (1)
The is_alive() operation returns False if the <<Task>> has died, i.e. it has exceeded its deadline by maximum delay. The <<Task>>s whose execution delay does not lead to critical outcomes can’t die. The estimate_duration() operation purpose is to determine the theoretical duration of a <<Task>> based on its properties. The duration is extended by the minimum task duration passed as an argument to the operation. Use cases for these operations are outlined in Table 1. The properties of the example <<Task>>s are as described:
  • effect – contains <<Task>>’s current position and state of <<Robot>> and <<Environment>> that the <<Task>> sets during execution or if finished. The examples are the <<Robot>>’s localisation or end-position of an object being transported,
  • deadline – the time during the <<Scenario>> by which we would like the <<Task>> to be completed,
  • priority – a priority of the <<Task>> (required by some <<DecAgent>>s), the higher, the more important the task is,
  • preemptive – defines if the <<Task>> can be interrupted after the <<Robot>> starts to work on it,
  • type – the name of the <<Task>> type in text form, useful for presenting the system’s state and scheduling algorithms assessments,
  • estimated_duration – estimated time required to complete the <<Task>> if its execution is started at the current time,
  • distance_from_robot – distance from the <<Robot>>’s position to the <<Task>>’s position,
  • deathtime – the time at which the task exceeds the deadline by maximum delay. Exceeding this time violates safety; thus, if deathtime is up, the <<Task>> is terminated, and the <<Scenario>> is failed.
These properties of <<Task>> are set and updated by the <<Scenario>> or its <<ScenarioPlugin>>s and used by the <<DecAgent>>s (to decide which <<Task>> should be performed) and <<EvalFunction>>s (to assess the current situation).
3.4 Benchmark configurability
Apart from <<Scenario>> and its <<Task>>s the <<TaBSASystem>> requires other blocks to function, such as <<DecAgent>> and its <<AgentPlugin>>s [UC1.3], <<EvalFunction>> [UC1.4], <<ScenarioPlugin>>s, and <<Robot>> [UC1.2.1]. These blocs depend on specific applications; thus, the users must configure them accordingly. They must implement the following operations and communication interfaces.
3.4.1 Configuring the Decision Agent
The <<DecAgent>> has one basic operation select_task() for selecting a <<Task>> for execution. To make this decision, it receives the list of jobs, the current time, and the output of the <<EvalFunction>> from the previous decision. The job list contains tasks already submitted but not completed (the ’jobs’ list of the <<Scenario>>). The list may contain more <<Task>>s than the <<DecAgent>> can process. If this is the case, it is at the <<DecAgent>>’s discretion to limit the number of <<Task>>s to be considered. The functionality of the agents can be extended using <<AgentPlugin>>s.
3.4.2 Configuring the Evaluation Function
The <<EvalFunction>> implements the calculate_results() operation. This function receives from the system the current time and the <<Task>> selected by the <<DecAgent>> for execution. The type of value returned by the operation is intentionally undefined so that it can be freely extended according to the user’s needs. For example, it could contain statistical data on the robot’s workflow or a quantitative assessment. The only requirement is to have a ’terminate’ boolean variable.
3.4.3 Creating the Agent Plugins
An <<AgentPlugin>> extends an agent with a selected capability of considerable complexity. Each <<AgentPlugin>> has its unique interface, depending on how it works. This is due to the infinite number of potential skills with which a <<DecAgent>> can be extended. <<DecAgent>>s wishing to use them must adapt to the way it works. By separating a capability as a <<AgentPlugin>>, there is no need to implement it individually for each <<DecAgent>>, and new <<DecAgent>>s are not forced to expand the code of the old ones to have it. Any <<DecAgent>> can use these <<AgentPlugin>>s. An illustrative example of an <<AgentPlugin>> could be a ’task request predictor’ that estimates the timestamps of future tasks.
3.4.4 Creating the Scenario Plugins
<<ScenarioPlugin>>s’ role is job processing, facilitating the modification of their parameters according to the <<Scenario>> and the <<Environment>> as user needs. For this reason, they require implementing a single update_job() operation (run within <<Scenario>>’s update_job() operation – Fig. 4) that modifies the submitted <<Task>> based on the current time within the <<Scenario>>. By design, changes made by <<ScenarioPlugin>> should be characterised by significant complexity but potentially not desirable on every run. Separating them allows them to be selectively activated for the current <<Scenario>>. <<ScenarioPlugin>>s may build up extra domain knowledge into the <<Task>> structure. An illustrative example could be a ’task duration estimator’ that calculates <<Task>>’s duration estimate so <<DecAgent>>s may use it.
3.4.5 Configuring the Robot
As the name suggests, the <<Robot>> represents the machine whose work is evaluated during the <<Scenario>>. Different <<Robot>> controllers may have different structures and functions. For this reason, we deliberately do not detail the <<Robot>>’s operations and properties. We only require it to have an execute_step() operation that takes the currently selected <<Task>> to be executed and the <<Environment>> as arguments.
3.5 Test-case configuration
3.5.1 The robot’s environment generation
The <<Environment>> in which the <<Robot>> works is <<Scenario>>’s property. The static obstacles emulating walls for the test session are generated using the recursive division method configured with the provided parameters (Fig. 5). First, all rooms, walls, and doors are generated considering the provided constraints. After that, the number of pieces of furniture is added to the map so as not to stand in front of any door. The number of furniture pieces is limited but not constant, as some rooms may be too small to place that much furniture, or the furniture just won’t generate properly in a finite number of tries. A set number of manipulation-intended items is placed on each piece of furniture. Ultimately, we generate starting and destination poses for human actors and define their behaviour.
Figure 5: Configurable environment parameters
The probability map is calculated based on the pedestrian spawning probabilities set during the environment configuration (Fig. 5/Pedestrians). Different probabilities can be set for the rooms, doors, and map entrance. The example environment map is presented in Fig. 6. All of this satisfy [UC2.3], specifically [UC2.3.1] and [UC2.3.2].
Figure 6: Sample randomly generated map with objects (blue), calculated probability distribution for pedestrian spawning (the darker green the bigger probability), and example footprints of the pedestrians (red) at the system runtime
3.5.2 Task generation
The <<Task>>s implemented in the <<Scenario>> are randomly generated using a chosen seed. This allows us to recreate the <<Task>> configurations while testing different <<DecAgent>>s or plugin configurations without saving all the <<Task>> information, as per [UC2.1]. While simulating the <<Robot>>’s work in a specific <<Environment>>, different seeds may represent different timeframes during the day or even certain days of the week, month, or year. For each task, we generate the time when it is given to the robot, the deadline, and the necessary environment positions – for example, the robot’s target position or positions of potential manipulation objects. The times must fit between the first step of the execution and the last step of the scenario. The positions are generated while considering the obstacles in the <<Environment>>. For example, the <<Robot>>’s position should be able to navigate to, and the object’s position should be within one of the existing pieces of furniture.
4 Example system
4.1 System setup
In the following section, we describe an example Mobile Manipulator System based on the <<TaBSASystem>>. It contains its own components based on the <<Scenario>>, <<ScenarioPlugin>>, <<Task>>, <<Robot>>, <<DecAgent>>, <<AgentPlugin>>, and <<EvalFunction>> stereotypes. They are graphically presented in Fig. 7.
Figure 7: Mobile Manipulator System
4.1.1 Tasks
Mobile Manipulator Task is based on the <<Task>> stereotype but implements numerous additional operations and properties (see Fig. 8).
Figure 8: Mobile Manipulator Task
The Fall <<Task>> is meant to imitate the situation in which the <<Robot>> is informed about an elderly falling to the ground. The <<Robot>> moves to the chosen location and waits there for a given time duration. This imitates a brief examination of the elderly health. This <<Task>> has the highest ’priority’, cannot be interrupted, and is terminated about 15 minutes after its ’deadline’. The Patrol <<Task>> resembles the <<Robot>> moving down the path from the starting point to the end. This <<Task>> can be interrupted at any time and cannot die. Patrol’s ’priority’ is determined randomly when it is created, but it can never be higher than the Fall’s ’priority’ value. Pick <<Task>> and Place <<Task>> are not used on their own. They are responsible for the <<Robot>> picking up and placing the object. As our objects are represented merely by the points on the map, the <<Task>>s are performed by standing for a set amount of time near the object. These <<Task>>s received a ’priority’ of 0 for the purposes of the Pick And Place <<Task>> - they aren’t used independently, so they don’t need a realistic ’priority’ value, and we didn’t want their priority to influence complex <<Task>>s.
To represent tasks composed of several simpler tasks, we prepared the Complex <<Task>>. This task possesses a list of ’subtasks’ that are worked on one after another. For this reason, the <<Task>>’s ’estimated_duration’ equals $\sum\limits_{subtasks}(subtask.estimated\_duration)$. The <<Task>>’s ’priority’ was experimentally set to $\max(subtask.priority:subtasks)$ to the sum of priorities of subtasks. The only composite <<Task>> prepared within the system was Pick And Place. It currently consists of 3 subtasks: Pick, Patrol, and Place. This <<Task>> aims to transport an item from one place to another. We set the Pick And Place <<Task>> as not interruptable. As these <<Task>>s utilize objects placed in the <<Environment>>, they tend to start and end close to each other.
4.1.2 Decision agents
We have implemented multiple <<DecAgent>>s to test the effectiveness of different <<Scenario>> execution approaches. Despite their inner workings, all <<DecAgent>>s respect the ’preemptive’ property of the <<Task>>s and won’t interrupt the <<Task>> that shouldn’t be paused during execution.
Simple-longest and Simple-shortest <<DecAgent>>s select the longest and the shortest <<Task>> for the execution, respectively. Additionally, they possess the ’hesitance’ property. Every time the <<Task>> is selected for execution, a random number between 0 and 1 is generated. If it is smaller than the hesitance value, the <<DecAgent>> returns the same <<Task>> as in the previous decision without checking anything (if the <<Task>> is still within the received ’jobs’).
The Distance <<DecAgent>> chooses the <<Task>> that is the furthest from the <<Robot>>. Additionally, the <<DecAgent>> possesses the ratio property used to offset the distance by a fraction of the <<Task>>’s estimated duration. If the ’ratio’ is non-zero, the <<DecAgent>> effectively chooses the <<Task>> with the highest ’dist_score’ calculated by subtracting the ’ratio’ times the <<Task>>’s ’estimated_duration’ from the <<Task>>’s ’distance_from_robot’:
$dist\_score=$ $\max_{jobs}(job.distance\_from\_robot-ratio*$ (2)
$*job.estimated\_duration)$
That means that a higher ratio causes the <<DecAgent>> to favour shorter <<Task>>s.
The Scheduler <<DecAgent>> uses the Request Table <<AgentPlugin>>. Request Table sorts the received <<Task>>s by highest ’priority’. <<Task>>s with the same ’priority’ are then sorted by the earliest ’request_time’. Then the <<AgentPlugin>> creates a schedule of sorts using <<Task>>s’ ’deadlines’ and ’estimated_durations’, returning the lists of scheduled and rejected <<Task>>s. Going through the sorted list, if the <<Task>>’s time from start to finish collides with any other <<Task>> in the scheduled list, the <<Task>> is assigned to the rejected list and vice versa. In the end, the Scheduler <<DecAgent>> receives both of the lists. It returns the <<Task>> from the scheduled list, which time of execution from start to finish contains the current time. If there is no such <<Task>>, the <<DecAgent>> doesn’t return any of them. The <<Task>>s which aren’t worked on at their scheduled time will not be completed in the free time.
The last <<DecAgent>> is called the DQN because it uses the neural network trained using the Deep Q Network [26] (DQN) algorithm to choose the <<Task>> for execution. The network’s input is of the shape [number of <<Task>> types][number of <<Task>>s per type][3] as presented in Fig. 9. The proposed system also contains the interface for the RLlib library [22]. The RLlib contains multiple Reinforcement Learning (RL) state-of-the-art methods that can also be tested for training and benchmarking algorithms scheduling tasks for mobile robots working in dynamic environments, so alternatively, many other RL algorithms can be used in this block. Considering that only Fall, Patrol, and Pick And Place <<Task>>s are directly used within the Service Scenario, in our case, the number of <<Task>> types is 3. For each <<Task>> type, the set number of <<Task>>s is selected by the earliest deadline. Sometimes, the amount of <<Task>> may be too low to fill the whole input. In such cases, the empty places are filled with zeros. For each <<Task>>, three properties are placed inside the network’s input: ’estimated_duration’, ’preemptive’ (represented as integer), and ’distance_from_robot’. <<Task>>s placed within the network’s input are remembered as a list. The network processes the input and outputs the values for each <<Task>>. The highest value marks the position of the <<Task>> to be executed within the saved list. The important process of training the DQN <<DecAgent>> (as per [UC2]) takes place inside its select_task() operation using deep reinforcement learning. After the <<DecAgent>> calculates the new input (shape presented in Fig. 9) it can supply it along with the previous input, selected <<Task>>, and ’reward’ from the <<EvalFunction>> to the training memory used to teach the network. Training the <<DecAgent>> is a long process and requires running multiple <<Scenario>>s using the same <<DecAgent>> (as presented in Fig. 3) as per [UC2.4].
Figure 9: DQNAgent’s network input and output
Apart from the Request Table, our system also contains another <<AgentPlugin>> – Task Predictor. Its purpose is to predict <<Task>>s that can be requested in the near future before they are called. This <<AgentPlugin>> possesses an instance of the Request Table <<AgentPlugin>>. Its most important part is a neural network with Long Short-Term Memory (LSTM) network [31] layers. The plugin is meant to predict <<Task>>s within the set time horizon divided into slots. For each slot, 3 values are calculated: number of <<Task>>s within the slot, number of <<Task>>s set to be executed within the slot (per information from Request Table), and the sum of the ’priorities’ of executable <<Task>>s within the slot. These values, along with the seed used for <<Task>> generation, are passed to the network. The network output consists of the same three values for the same amount of slots but for time horizon counted since one slot in the future. The visualisation of this network’s input and output can be found in Fig. 10. Calculating the differences between the output and the input allows predicting if new <<Task>>s will be added within these time slots in the future.
Figure 10: Task Predictor’s network input and output
4.1.3 Scenario configurations
The example Service Scenario, based on the <<Scenario>> stereotype, has a Mobile Manipulator <<Robot>>, a single <<ScenarioPlugin>> – Duration Estimator Network, and only Mobile Manipulator Task <<Task>>s (Fig. 11). The <<Scenario>> is meant to represent several hours of work for the mobile service <<Robot>>. The <<Robot>> has a set velocity for traveling between <<Task>>s. A certain number of <<Task>>s of each type – Fall, Patrol, and Pick And Place – is generated within these hours. The properties of these <<Task>>s are updated every time by the update_job() operation: estimate_duration() is called, ’distance_from_robot’ is calculated, and every path that needs to be planned within the <<Task>> is refreshed using the current state of the environment.
Figure 11: Service Scenario
Duration Estimator Network <<ScenarioPlugin>> is as the name suggests meant to estimate the duration of the <<Task>>. Despite every <<Task>> in our Mobile Manipulator System having its own estimate, in the real world, these numbers may be offset by an inconsistent environment caused by, for example, humans walking around or misplacement of furniture and objects. Therefore precise prediction of the <<Task>>’s duration is impossible. The plugin consists of a neural network trained offline on data gathered from multiple runs. To mitigate overtraining the network, during data gathering, ’durations’ and ’deadlines’ of <<Task>>s are slightly modified by adding noise and additional <<Task>>s are added randomly to the <<Scenario>>. The network has a single output representing the predicted <<Task>> duration. Its input, on the other hand, consists of the following values: the day when the predicted <<Task>> is performed (seed from <<Task>> configuration), current time, <<Task>>’s ’deadline’, <<Task>>’s current position, <<Task>>’s goal position (if it requires movement, otherwise current position again), <<Task>>’s current ’distance_from_robot’ and <<Task>>’s ’priority’. The structure is presented in Fig. 12.
Figure 12: Duration Estimator Network’s network input and output
4.1.4 Evaluation functions
Two <<EvalFunction>>s are currently implemented within the Mobile Manipulator System: DQN Eval and Statistic Eval. Both of them have their specific purposes.
DQN Eval is the <<EvalFunction>> created mainly to train the DQN <<DecAgent>> because the <<DecAgent>>’s most important part is the neural network trained using reinforcement learning. It requires a proper reward function to learn how it should operate. Therefore, DQN Eval functions as both an <<EvalFunction>> for the <<TaBSASystem>> and a training reward for DQN <<DecAgent>>. DQN Eval returns a ’terminate’ flag if all <<Task>>s are completed or one of the <<Task>>s is terminated. It also returns the ’reward’ parameter that depends on the <<DecAgent>>’s performance. If the <<DecAgent>> chooses an existing <<Task>>, one or all <<Task>>s are completed, the reward is positive. If the <<Task>> chosen by the <<DecAgent>> doesn’t exist or one of the <<Task>>s is terminated, the reward takes the form of a negative penalty. The <<DecAgent>> is also slightly penalized for switching <<Task>>s they work on to discourage it from jumping between <<Task>>s too often. All of these values are represented by appropriately named parameters. The reward function is presented in the equation (3). The additional parameter— ’penalty_change_job’ is added to the reward only in the last two cases if the current action is different than the ’previous_action’ property.
$\footnotesize reward=\left\{\begin{array}[]{ll}reward\_all\_complete&if\,all\,tasks\,are\\ &\quad completed\\ penalty\_dead\_job&if\,at\,least\,one\,job\\ &\quad is\,dead\\ reward\_job\_complete&if\,job\,was\,just\\ &\quad completed\\ 0&if\,there\,is\,no\,jobs\\ &\quad in\,the\,list\\ reward\_real\_job&if\,real\,job\,was\\ &\quad selected\\ penalty\_nonexistent\_job&if\,nonexistent\,job\\ &\quad was\,chosen\\ \end{array}\right.$ (3)
Statistic Eval aims to numerically assess the work of any <<DecAgent>> it monitors using statistics. This makes it good for every <<DecAgent>> in the <<TaBSASystem>>. It allows the user to verify what the decisions made by the <<DecAgent>> accomplish in the chosen <<Scenario>>. Statistic Eval returns the ’terminate’ flag if all <<Task>>s are completed, one of the <<Task>>s is terminated or the <<DecAgent>> switches between <<Task>>s and returns to the same one for the third time within 3 minutes. Other useful values returned by the evaluation function on each iteration are represented by the following properties:
  • full_travel_distance – the full travel distance of the <<Robot>> calculated by comparing the <<Robot>>’s current pose to the remembered ’last_robot_pos’,
  • num_of_tasks_completed – the number of completed <<Task>>s of every type in ’task_types’,
  • task_completion_to_deadline – the difference between the <<Task>>’s completion time and the original ’deadline’,
  • task_completion_to_deathtime – the difference between the <<Task>>’s completion time and potential time of termination, None if a <<Task>> can’t be terminated,
  • num_of_tasks_interrupted – number of <<Task>>s’ interruptions per <<Task>> type in ’task_types’,
  • task_interruptions – number of <<Task>>s’ interruptions for each <<Task>> in <<Scenario>>’s ’task’ list,
  • num_of_human_abandonement – number of <<Task>>s’ interruptions for each <<Task>> and number of times the <<Robot>> has abandoned a human in need.
The last statistic is specific to our system and represents the number of situations where the <<Robot>> drove within ’abandonment_distance’ of any Fall <<Task>> without working on it.
4.2 Example benchmarking sessions
4.2.1 Decision agent training
As mentioned before, DQN <<DecAgent>> requires a trained neural network to operate. Training this network may be considered an example of a successful benchmarking session utilising our Mobile Manipulator System. The <<DecAgent>> was taught using reinforcement learning (DQN algorithm). The training consisted of the <<DecAgent>> rehearsing numerous <<Scenario>>s. During the training, the <<Scenario>>s reflected 4 hours of social <<Robot>> work, during which the <<DecAgent>> was expected to complete 12 <<Task>>s of each of the 3 types. In each, the <<Task>>s were generated using a new seed. The DQN <<DecAgent>>’s actions were evaluated using DQN Eval, and the reward received during this stage was used as a reward for reinforcement learning. Learning took several million steps, each representing 5 seconds of virtual time. Individual <<Scenario>>s ended if all <<Task>>s were completed, one of the <<Task>>s died or <<Task>> completion exceeded the <<Scenario>> time. Once training was complete, the network was saved for further use. During the project, the training was repeated several times to verify differences in the performance of the different network structures and the value of the reward for individual actions.
4.2.2 Agent comparison
Another example of a benchmarking session could be an attempt to compare the performance of individual decision-making <<DecAgent>>s. Such an exercise makes it possible to determine the best one and to identify potential errors in its performance. Initially, 50 <<Scenario>>s with random <<Task>>s were generated for each <<DecAgent>> under test. The <<Scenario>>s lasted 4 hours and contained 12 <<Task>>s of each of the three types. This time the Statistic Eval was used to evaluate the performance of the <<DecAgent>>s. The results of each run were saved for analysis. Some of the results for six different <<DecAgent>>s (Distance with ’ratio’ 0.5, two versions of DQN, Scheduler, Simple-longestwith ’hesitance’ 0.5, and Simple-shortest with ’hesitance’ 0.5) are shown below, as they presented us with useful information about the <<DecAgent>>s and allowed us to improve them.
Fig. 13 includes statistics for scenario termination cause for different types of <<DecAgent>>s presented as bars. We can see that Simple-shortest and Distance <<DecAgent>>s were moderately successful in completing more than half of the <<Scenario>>s. Scheduler and Simple-longest <<DecAgent>>s both failed for different reasons. If two Fall <<Task>>s were to be scheduled at the same time, Scheduler would only complete one of them. Simple-longest agent, on the other hand, prioritizes long <<Task>>s, which causes him to start with Pick And Place <<Task>>s, which tend to be longer than the others. DQN <<DecAgent>> didn’t have a good outcome during this test either, but something else was also spotted. After training the first DQN <<DecAgent>>, a lot of <<Scenario>>s ended due to running out of time. This fact prompted us to reevaluate the code and discover that the penalty for doing nothing was indeed wrongly applied as a positive number. This was fixed for the second network, which, as one can see, doesn’t display the same behaviour.
Figure 13: Termination statistics for 6 <<DecAgent>>s (Distance, two versions of DQN, Scheduler, Simple, and Simple-shortest)
Fig. 14 shows the absolute difference between <<Task>>s’ completion times and their original ’deadlines’ sorted by <<Task>> type. As can be seen, one of the <<DecAgent>>s seems to be much better than the others - Scheduler - as it displayed much smaller disparities. This means that <<Task>>s worked on by this agent were completed close to the original ’deadlines’, which may be a desirable behaviour in some cases. All of the other <<DecAgent>>s start performing <<Task>>s as fast as they come, which causes the differences to be quite large. We can also observe that the biggest differences are visible for Pick And Place <<Task>>s, which suggests that they are usually the longest to complete or are susceptible to moving pedestrians.
Figure 14: Difference between real completion time and the ’deadline’ for different <<Task>> types for different <<DecAgent>>s
Fig. 15 shows the number of <<Task>>s of each type completed by each <<DecAgent>> during the mentioned 50 <<Scenario>>. These plots also confirm that most <<DecAgent>>s work as intended despite failing the <<DecAgent>>s. For example, the Scheduler has more Fall <<Task>>s completed than other types, as they have the highest manual ’priority’. On the other hand, Simple-longest chooses to work on the longest <<Task>>s first, so it prioritizes Pick And Place <<Task>>s. The biggest surprise is the first DQN, which seems to have completed almost as many fall <<Task>>s as the Distance <<DecAgent>> but didn’t complete any <<Scenario>>s. This further confirms the need to retrain the <<DecAgent>> with different parameters.
Figure 15: Number of completed <<Task>>s per type for different <<DecAgent>>s
After that, another experiment was performed, in which each <<DecAgent>> completed the same 100 <<Scenario>>s (same seed used for <<Task>> generation). This allowed us to inspect the differences in <<DecAgent>>s’ behaviours closely. Figures 16 and 17 show how the same scenario was completed by Simple-shortest and Distance <<DecAgent>>s. The horizontal axis represents the passage of time, while the currently performed <<Task>> is identified by its colour and order in the <<Task>> list of the <<Scenario>>. For example, a green line at 6 represents the sixth generated Pick And Place task. Comparing the plots, we can see that the <<Task>>s were completed in a slightly different order at the beginning due to differences in decision-making. On the other hand, since about 1700th second, the outcomes are almost identical, probably because these <<Task>>s were given to the agents at the same time late into the day.
Figure 16: Scenario course as performed by Simple-shortest <<DecAgent>>
Figure 17: Scenario course as performed by Distance <<DecAgent>>
5 Summary
This article tackles the problem of assessing task-scheduling algorithms in the mobile robot domain. We propose a benchmarking framework that encapsulates the scheduling algorithm with the <<DecAgent>> interface. Numerous algorithms are optimised for given robot applications, robot types, or environments. Therefore, our framework allows the configuration of the robot’s kinematics, tasks, environment, and testing scenarios. By the configuration, the framework reflects the desired robot system and its environment setup. The system is open-source, and its repository is shared [^2] https://github.com/RCPRG-ros-pkg/Smit-Sim. Therefore, our experiments can be easily re-executed, and the framework can be used to evaluate new or well-known algorithms in the domain of mobile robots.
The framework incorporates the generators for schedules of <<Task>> requests, trajectories of moving humans, and maps with objects for manipulation and rooms. These generators can be used in training task-scheduling algorithms and for statistic-based evaluation of these or other algorithms.
We have validated the framework by configuring it for a mobile service <<Robot>> application, teaching AI-based algorithms using DQN, and assessing it against basic scheduling algorithms. As a result of the framework application, we could identify flaws in both coding and decision-making. Ultimately, we could understand and graphically present the core differences between the selected algorithms at chosen configurations. In the given test case, scheduling algorithms that chose the shortest (Simple-shortest <<DecAgent>> with ’hesitance’ 0.5) or the furthest (Distance <<DecAgent>> with ’ratio’ 0.5) task managed to complete the <<Scenario>>s respectively  90% and  60% of the time. They also didn’t seem to favour any specific type of <<Task>>. The algorithm choosing the longest <<Task>> (Simple-longest<<DecAgent>> with ’hesitance’ 0.5) did the worst (with all <<Scenario>>s fatally failed), followed closely by the AI-based ones (both versions of DQN <<DecAgent>>) and the one based on the strict schedule (Scheduler <<DecAgent>>). In the first case, the <<DecAgent>> favours executing Pick And Place, while the others Fall <<Task>>s. Of all the algorithms, only the schedule-based one completed the <<Task>>s almost at the set ’deadline’.
From testing the algorithms in the proposed benchmarking system, we not only select the best algorithm but also learn about the <<Robot>> application. During the investigation of our example, we learned that both of the best algorithms managed to get ahead by completing as many <<Task>>s as fast as possible by favouring shorter <<Task>>s. Additionally, the Distance <<DecAgent>> avoided tasks that are close to each other. Pick And Place <<Task>>s tend to start and end close to each other and are often the longest to perform, so algorithms that favour them for any reason are likely to fail. Seeing that most scenarios ended due to the "dead" status (which happens when Fall <<Task>> is not completed for some time after its ’deadline’), we find that keeping deadlines is also critical in this application. Thus, the scheduling algorithms that are aware of deadlines should perform better than the ones that are not. On the other hand, a <<DecAgent>> may concentrate on Fall <<Task>>s but can still fail if it uses a subpar scheduling algorithm.
Scheduling tasks of mobile robots is challenging in a dynamic environment. However, based on the experience gained in this research, we find the advanced scheduling of manipulation tasks the next valuable target. Advanced scheduling must involve semantic planning of the tasks, which is a state-of-the-art solution to managing complexity in robot tasks [29].
[…]
Acknowledgements
The research was funded by the Centre for Priority Research Area Artificial Intelligence and Robotics of Warsaw University of Technology, Poland, within the Excellence Initiative: Research University (IDUB) programme, agreement no. 1820/336/Z01/POB2/2021. The work of Kamil Młodzikowski and Dominik Belter was supported by the National Science Centre, Poland, under research project no UMO-2023/51/B/ST6/01646.

### v3

@@
-
+TaBSA – A framework for training and benchmarking algorithms scheduling tasks for mobile robots working in dynamic environments
 Abstract
-Service robots work in a changing environment habited by exogenous agents like humans.
-In the service robotics domain, lots of uncertainties result from exogenous actions and inaccurate localisation of objects and the robot itself.
-This makes the robot task scheduling problem challenging.
-In this article, we propose a benchmarking framework for systematically assessing the performance of algorithms scheduling robot tasks.
-The robot environment incorporates a map of the room, furniture, transportable objects, and moving humans.
-The framework defines interfaces for the algorithms, tasks to be executed, and evaluation methods.
-The system consists of several tools, easing testing scenario generation for training AI-based scheduling algorithms and statistical testing.
-For benchmarking purposes, a set of scenarios is chosen, and the performance of several scheduling algorithms is assessed.
-The system source is published to serve the community for tuning and comparable assessment of robot task scheduling algorithms for service robots.
-The framework is validated by assessment of scheduling algorithms for the mobile robot executing patrol, human fall assistance and simplified pick and place tasks.
+This article introduces a software framework for benchmarking robot task scheduling algorithms in dynamic and uncertain service environments.
+The system provides standardized interfaces, configurable scenarios with movable objects, human agents, tools for automated test generation, and performance evaluation.
+It supports both classical and AI-based methods, enabling repeatable, comparable assessments across diverse tasks and configurations.
+The framework facilitates diagnosis of algorithm behavior, identification of implementation flaws, and selection or tuning of strategies for specific applications.
+It includes a SysML-based domain-specific language for structured scenario modeling and integrates with the ROS-based system for runtime execution.
+Validated on patrol, fall assistance, and pick-and-place tasks, the open-source framework is suited for researchers and integrators developing and testing scheduling algorithms under real-world-inspired conditions.
 keywords:
-[…]
-{graphicalabstract}[Uncaptioned image]
-{highlights}
-Goal: Comparable evaluation of various algorithms scheduling diverse robot tasks.
-Method: Kinematic simulation executed in configured or generated scenarios and environments.
-Result: The framework teaches an RL-based algorithm, and other algorithms are compared.
-Conclusion: The comparison shows algorithm performance for specific scenarios and tasks.
-1 Introduction
+robot scheduling algorithms , benchmarking tools , dynamic environments , reinforcement learning
+journal: SoftwareX
+Metadata
+Nr. Code metadata description Please fill in this column
+C1 Current code version 1.0.0
+C2 Permanent link to code/repository used for this code version https://github.com/RCPRG-ros-pkg/Smit-Sim/tree/v1.0.0
+C3 Permanent link to Reproducible Capsule https://github.com/RCPRG-ros-pkg/Smit-Sim/blob/v1.0.0/Dockerfile
+C4 Legal Code License MIT License
+C5 Code versioning system used git
+C6 Software code languages, tools, and services used python3.7, bash, ROS melodic
+C7 Compilation requirements, operating environments & dependencies https://github.com/RCPRG-ros-pkg/Smit-Sim/blob/main/requirements.txt
+C8 If available Link to developer documentation/manual https://github.com/RCPRG-ros-pkg/Smit-Sim/blob/main/README.md
+C9 Support email for questions daniel.gieldowski@pw.edu.pl
+Table 1: Code metadata (mandatory)
+1 Motivation and significance
 1.1 Problem statement
-Humans have honed the ability to manage and prioritise multiple tasks for centuries.
-Today, the challenge lies in transferring this capability to robots.
-This task has become increasingly important due to the growing need for robotisation in various domains facing staff shortages.
-Thus, robots ability to multitask efficiently becomes critical, and the proposed scheduling algorithms should be assessed systematically.
-Solving the problem of robot task management requires adapting task scheduling solutions to real-world environments characterised by dynamic changes [8], uncertainty [13], and interactions with exogenous actors such as humans [12, 18].
-This presents a significant challenge, as changing environments affect the schedules and action plans required to perform the requested tasks.
-For safety and effectiveness reasons, task scheduling algorithms should be rigorously tested in simulated environments [3, 16].
-Evaluations of these algorithms may involve statistical analysis or proving their optimality.
-The latter is especially difficult in changing, inhabited environments, particularly for neural-network-based scheduling algorithms [15].
-Task scheduling for mobile robots in dynamic environments has been extensively studied, with researchers exploring both traditional and AI-based methods to address the inherent challenges.
-Classical approaches often rely on heuristic or optimization-based strategies.
-However, these methods struggle with scalability and adaptability in environments with frequent changes.
-Modern AI-driven methods, such as reinforcement learning, have shown promise in addressing dynamic scheduling.
-For example, [30] review studies focused on Reinforcement Learning techniques that have been used for dynamic task scheduling.
-[33] argue that a Hybrid Q-learning approach allows for maintaining optimal scheduling.
-Furthermore, they state their reinforcement learning method is resilient to non-deterministic environments.
-They validate the algorithm in a robotic arm environment, and in the future, they want to focus on motion planning, collision avoidance, and multi-agent algorithms.
-Similarly, [36] proposed a deep multi-agent reinforcement learning approach to optimize task scheduling in human-robot collaboration, showcasing its effectiveness in dynamic environments.
-Yet, these techniques often require extensive training data and may falter in highly unpredictable scenarios.
-Simulation tools like Gazebo and frameworks such as ROSPlan [5] have facilitated the testing of scheduling algorithms, but they are limited in their ability to capture the complexities of human-robot interactions and real-world uncertainties.
-Despite these advancements, benchmarking service robot scheduling algorithms in such environments remains an open problem, as existing tools and methods fall short in addressing the unpredictable and stochastic nature of real-world settings, and transparent comparison of the methods is difficult.
-Thus, a framework for comparable benchmarking of the algorithms is mandatory.
-Although current frameworks and benchmarking systems provide essential tools for evaluating task scheduling algorithms, they often fall short in addressing the complexities of dynamic and uncertain environments in the field of service robotics.
-The framework introduced in this work aims to bridge this gap by offering a structured approach to training and benchmarking scheduling algorithms.
-The proposed framework explicitly accounts for real-world uncertainties and facilitates fair performance comparisons across diverse scenarios.
-The framework is adaptable, supports various environments and algorithms, and allows for systematic and repeatable testing against configurable scenarios.
-Importantly, the framework ensures that performance scores are comparable between different algorithms and scenarios.
-In addition, it provides tools for generating test cases and scenarios, which can be used to train AI-based scheduling algorithms.
-Additionally, this article introduces a SysML-based [35] domain-specific language (DSL) designed to support the presentation, documentation, and analysis of benchmarking systems developed with the proposed system.
-The DSL follows the Model-based Systems Engineering (MBSE) method used broadly for systems analysis, design verification and validation in the V-model [6] workflow.
+The growing demand for robotisation in sectors facing labour shortages highlights the need for autonomous robots capable of managing multiple tasks efficiently.
+A critical challenge in this context is designing and evaluating scheduling algorithms that can operate reliably in dynamic [1], uncertain [2], and human-inhabited environments [3, 4].
+While AI-based approaches [5, 6], such as reinforcement learning, offer promising solutions, their real-world deployment remains difficult due to the unpredictability of operational contexts and the lack of reliable, comparable evaluation tools.
+Existing simulation platforms and benchmarking frameworks, although useful, often fall short in capturing the complexity of real-world conditions and ensuring transparent comparison of algorithmic performance.
+As a result, the assessment of scheduling algorithms lacks consistency and reproducibility, limiting their integration into practical robotic systems.
+The goal of this work is to provide a software framework that enables systematic and reproducible evaluation of both existing and emerging scheduling approaches, offering a solid foundation for the development and comparison of new algorithms.
+Demonstrating strong performance in some scenarios or failure in others are both valuable outcomes, as they highlight the framework’s ability to reveal algorithmic strengths and weaknesses in dynamic and uncertain environments.
+To facilitate adoption, we provide a Dockerfile for easy deployment and an introductory video [^1] https://vimeo.com/1122196556 .
+1.2 Software capabilities
+The software framework is dedicated to training and benchmarking robot scheduling algorithms.
+It enables systematic, repeatable testing in configurable scenarios and ensures fair performance comparisons across algorithms.
+The framework supports uncertainty modeling, task-specific scenario generation, and comparative evaluation, making it suitable for both algorithm developers and mobile manipulator integrators.
+It also includes a Domain-Specific Language (DSL) based on Systems Modeling Language (SysML), which facilitates structured documentation and analysis of benchmarking setups using a model-based systems engineering approach.
+Additionally, the SysML-based description supports assessing software consistency between simulation and physical deployment via SPSysML [7].
+This article presents the framework’s architecture, usage, and validation.
+It also outlines its integration with robot control systems, including the ROS-based TaskER platform [8] for safe task switching.
+The proposed framework aims to standardize and accelerate the development of robust, multitasking robots for real-world applications.
 The framework’s capabilities are illustrated through the use case diagram Fig.
 1.
-As target users of the framework, we recognise the scientists developing scheduling algorithms and mobile manipulator integrators.
-The algorithm developers can use the framework to test their solutions against other algorithms and previous versions of their work in a comparable environment.
-The mobile manipulator integrators can mimic a use-case-specific environment in the framework and validate which scheduling algorithms best suit the tasks and environment of the robot application.
 Figure 1: Use cases of the proposed benchmarking system
-The framework should ease analysing how the scenario’s change influences the algorithm’s score.
-This article investigates how such a system should be structured, how it works, and how to use it to receive comparable results of different scheduling algorithms evaluated in various scenarios.
-The best-assessed scheduling algorithm must be integrated into the end product— robot control system.
-For the architecture-level integration, we have proposed the Simulation-Physical Modeling Language [9], and for the implementation-level integration, we have published the ROS [^1] Robot Operating System [19, 34]-based TaskER framework [10] that models the structure and behaviour of a multitasking robot.
-TaskER manages safe interruptions between different task contexts, allowing the robot to conduct safety-critical actions of the interrupted task before executing the switch to the next task.
-The contribution of this article can be summarized as follows:
-  • a comprehensive framework for training and benchmarking robot task scheduling algorithms, explicitly addressing real-world uncertainties in dynamic and human-inhabited environments.
-It supports systematic, repeatable testing across configurable scenarios while ensuring comparable performance evaluation.
-  • a dedicated domain-specific language for presenting, documenting, and analyzing benchmarking systems, enhancing transparency and standardization in the development and assessment of scheduling algorithms.
-  • an Open-Source and Practical Validation: The framework, validated through scenarios involving patrol, human fall assistance, and pick-and-place tasks, is published as an open-source tool, promoting community collaboration and the development of AI-based scheduling algorithms.
-In the next section 2, we summarise the state-of-the-art solutions to the problem of benchmarking task scheduling algorithms.
-In the following sections, we introduce the architecture of the proposed benchmarking system 3.1, the scope of possible benchmarking scenarios 3.2, the configuration of the system for testing specific algorithms and robots 3.4, and finally, we describe the developed tools for generating test-cases 3.5.
-In section 4, we describe the configuration of the example system that evaluates basic and neural-network-based algorithms.
-The neural network algorithms were taught on the test cases generated by the proposed tools.
-Finally, the evaluation results are presented and discussed in 4.2.
-2 Related work
-2.1 Simulators for developing and benchmarking in robotics
-Simulators play a crucial role in developing and benchmarking algorithms in the robotics domain.
-The Stage simulator, as discussed by [14], provides a simplified platform for benchmarking navigation systems in static environments.
-However, its project has been inactive for a long time, making its utility limited by the challenges of integrating recent algorithms.
-Additionally, the review of physics-based simulators by [23] emphasizes the importance of simulation accuracy in testing robotic controllers and the potential benefits of simplifying simulations for high-level task scheduling problems.
-The varying performance of simulators across different robotic domains, as noted by [7], further underscores the need for tailored benchmarking systems that account for the specific physical effects relevant to each application.
-Simulation is a crucial development tool for robotics, which allows avoiding challenges of learning on the real robot, especially when learning-based techniques are applied [17].
-This approach avoids the time-consuming experiments on real platforms and mitigates the risk of damaging the real robots.
-Moreover, the utilization of parallel architecture of the Graphical Processing Units enables simulating many robots at the same time and significantly reduces the training time [28].
-The new PhysX-based Orbit platform standardizes the benchmarking manipulation and robot locomotion tasks [25].
-2.2 Benchmarking task scheduling
-The problem of robot task scheduling in dynamic and uncertain environments has garnered significant attention in the field of robotics.
-In the general domain of task scheduling, several frameworks and benchmarking systems have been proposed.
-One notable framework is introduced by [27], which leverages large language models (LLMs) for task division and tool selection.
-The framework evaluates the performance of three LLMs by measuring the success rate and the order of tool usage in completing subtasks.
-This approach demonstrates the potential of LLMs in enhancing task scheduling through intelligent tool selection, although its primary focus is on utilising language models rather than the broader context of robot task scheduling.
-LLMs are also used to interpret vocal requests of tasks and identify their parameters [32].
-Another significant contribution is TASKOGRAPHY [1], a large-scale benchmark designed to evaluate robotic task planning over 3D scene graphs (3DSGs).
-This work also introduces SCRUB, a planner-agnostic strategy for adapting 3DSGs to improve planning performance, and SEEK, a procedure that enhances learning-based planners’ ability to exploit 3DSGs.
-These tools are particularly relevant for environments requiring detailed scene descriptions, although they focus more on planning than scheduling tasks in a dynamic environment.
-The survey by [4] provides a comprehensive overview of various task scheduling metrics, offering a foundational understanding of the criteria used to assess scheduling algorithms.
-This work, along with the review of metrics and benchmarking for parallel job scheduling by [11], highlights the diversity of metrics applicable to scheduling, though these works primarily concentrate on static environments.
-In the domain of task and motion planning (TAMP), [21] proposes a platform-independent evaluation method, which is crucial for assessing the generalizability of scheduling algorithms across different robotic platforms.
-Similarly, the work by [27] offers a structured framework for LLM-based AI agents, focusing on the execution of inference processes, but again, with a specific focus on language models rather than broader task scheduling in robotics.
-For benchmarking dynamic job scheduling, [24] presents the Dynamic Job Scheduler Benchmark (DJSB), a tool for comparing performance metrics across different scenarios.
-This tool, which allows for testing various resource management strategies, is particularly relevant for applications where resource availability and task demands change over time.
-The framework proposed by [2] addresses the PET-aware task-to-core scheduling problem, offering evaluation procedures and benchmarks for comparing scheduling algorithms.
-This work, along with the graph-based scheduling algorithm benchmarking by [20], contributes valuable methodologies for assessing the efficiency and performance of scheduling algorithms, particularly regarding resource utilization and execution time.
-In summary, while existing frameworks and benchmarking systems offer valuable tools and methodologies for evaluating task scheduling algorithms, there remains a need for a comprehensive system that addresses the unique challenges posed by dynamic and uncertain environments in the service robotics domain.
-Our proposed framework seeks to fill this gap by providing a systematic approach to training and benchmarking scheduling algorithms, incorporating real-world uncertainties and enabling fair comparisons across different scenarios.
-Besides the above core functional contribution, this article provides a domain-specification language (DSL) for the presentation, documentation, and analysis of the benchmarking systems that are built using the proposed framework.
-The DSL is expressed in Systems Modeling Language (SysML) [35].
-3 The framework for benchmarking and training scheduling algorithms
-3.1 The architecture
-TaBSA— the framework for Training and Benchmarking Scheduling Algorithms is a metamodel that, by configuration, gives a specific benchmarking system named TaBSA System.
+The proposed framework facilitates benchmarking and development of task scheduling algorithms in realistic robotic environments. It has been validated in three service robot scenarios:
+  • 1.
+    patrol and monitoring tasks in dynamic spaces,
+  • 2.
+    assistance in human fall detection and intervention,
+  • 3.
+    pick-and-place operations involving object manipulation.
+We provide ready-to-use scenario generators, benchmarking tools, and an interface for training AI-based algorithms under controlled uncertainty.
+Researchers and integrators can test algorithms in reproducible settings and evaluate their robustness to changing conditions.
+The main contribution of this work can be summarized as follows:
+  • 1.
+    a complete framework for training and benchmarking robot task scheduling algorithms in uncertain, human-inhabited environments,
+  • 2.
+    integration of a domain-specific language for formal documentation and analysis of benchmarking systems,
+  • 3.
+    open-source implementation validated in real-world-inspired service robot scenarios, enabling repeatable and comparable evaluation of scheduling methods.
+2 Software description
+TaBSA— the framework for Training and Benchmarking Scheduling Algorithms bases on a metamodel that, by application, gives a specific benchmarking system named TaBSA System.
 The TaBSA System is organised in a way presented in the SysML block definition diagram in Fig.
-2.
-The diagram is the part of the metamodel for the variety of particular TaBSA System configurations (realisations).
-For each block, a dedicated stereotype was introduced to simplify the specification of these realisations.
-The TaBSA System addresses the use cases [UCX] from Fig.
-1, by, a.o., reflecting the entities from these use cases as dedicated stereotypes: <<TaBSASystem>> – [UC1] and [UC2], <<Scenario>> – [UC1.2], <<Robot>> – [UC1.2.1], <<Task>> – [UC1.2.2], <<Environment>> – [UC1.2.3], <<DecAgent>> – [UC1.3], <<EvalFunction>> – [UC1.4].
-Figure 2: TaBSA system structure
+2(a).
+(a) TaBSA system structure
+(b) TaBSA system main operation
+Figure 2: TaBSA structure and behaviour
 All system elements that can constitute the variety of configurations are composed (filled rhombus) into <<TaBSASystem>>.
@@
 The <<TaBSASystem>> general behaviour is also a part of the metamodel.
-It is depicted in activity diagram 3.
+It is depicted in activity diagram 2(b).
 It provides a scheme for repeatable execution of configurable scenarios [UC1].
@@
 A detailed description of the particular operations mentioned in the diagram is presented in the following part of the article.
-Figure 3: TaBSA system main operation
-3.2 Scope of benchmark scenarios
+2.1 Framework configurability
+2.1.1 Scope of benchmark scenarios
 Apart from using a specific <<Robot>>’s model, the <<Scenario>> defines important aspects necessary for testing and training task management agents.
@@
 The scenario’s most important operations are presented in the form of pseudocode in Fig.
-4.
+3.
 Figure 3: Pseudocodes for some operations of Scenario, Task, and EvalFunction
@@
 We can define the different types of <<Task>>s our robot is supposed to perform as long as they can be represented in our restricted environment, as per [UC1.2.2].
-This includes, a.o., simple operations such as navigation, picking, and placing of objects.
-Once we decide, we can fill the list with the selected amount for each type of <<Task>>.
-Each <<Task>> must possess the ’request_time’ property that represents at what time it is presented to the robot.
-The <<Task>>’s operations include is_called() and is_completed().
-Apart from ’tasks’, the scenario also possesses the ’jobs’ list.
-This list contains solely the <<Task>>s that have already been called (is_called() operation returns True), so it can never contain a <<Task>> that can’t be found on the ’tasks’ list.
-The <<Scenario>> defines the ’environment’ (of <<Environment>> stereotype) consisting of a flat map with walls, door openings, and furniture (simple shapes blocking the movement), items placed on the furniture, and moving humans represented by footprints of their legs, as per use case [UC1.2.3].
-Simulated humans move on the map using simple kinematic equations and PRM algorithm for path planning, and their behaviour defines what happens after they reach the end of their path.
-They can circle between two points, choose a new destination, instantly change their position and choose a new destination, or just disappear.
-Humans are dynamically embedded on the map to make it possible for the robot to consider them during navigation.
-This represents the system incorporating external sensors for human detection.
-Manipulation targets are described solely by objects’ location.
+The <<Scenario>> defines the <<Environment>> consisting of a flat map with walls, door openings, and furniture (simple shapes blocking the movement), items placed on the furniture, and moving humans represented by footprints of their legs, as per use case [UC1.2.3].
+Simulated humans move on the map using simple kinematic equations and PRM algorithm for path planning, and they execute complex behaviours like going though list of destinations.
+Humans are embedded on the map, so the robot has to avoid them.
+Robot manipulation targets placed in the environment are moveable by the robot.
 These provide a highly simplified representation of a real robot’s environment but are sufficient to evaluate <<Task>> execution order and allow training of neural network-based algorithm in a reasonable time.
 Due to its simplicity, it also satisfies the use case [UC2.2].
-Table 1: Task operations as called by different components of our example benchmarking system — Mobile Manipulator System (for example pseudocodes see Fig. 4)
-Operation Called by
-is_called() <<Scenario>> - to determine when to add the <<Task>> to ’jobs’ list
-is_completed() <<Scenario>> - to determine when to remove <<Task>> from ’jobs’ list
-DQN Eval (Fig. 7) - to determine when the <<Task>> is completed and proper reward should be granted
-Statistic Eval (Fig. 7) - to determine when the <<Task>> is completed to calculate time difference between
-     planned and final completion
-is_alive() DQN Eval (Fig. 7) - to determine if the <<Task>> died and therefore <<Scenario>> should terminate
-Statistic Eval (Fig. 7) - to determine if the <<Task>> died and therefore <<Scenario>> should terminate
-work_for() Mobile Manipulator (Fig. 7) - inside it’s execute_step() operation, it is passed as argument by the
-     <<Scenario>>’s work() operation
-estimate_duration() <<Scenario>> - to update <<Task>>’s estimated durationbased on it’s own properties. It is called
-     by update_job() operation
-3.3 Task features
+2.2 Tasks adaptation
 Among the operations of <<Task>>, the most important is work_for(), whose implementation determines how the <<Task>> is executed.
@@
 Deathtime equals:
-$deathtime=deadline+max\_delay.$ (1)
-The is_alive() operation returns False if the <<Task>> has died, i.e.
-it has exceeded its deadline by maximum delay.
-The <<Task>>s whose execution delay does not lead to critical outcomes can’t die.
-The estimate_duration() operation purpose is to determine the theoretical duration of a <<Task>> based on its properties.
-The duration is extended by the minimum task duration passed as an argument to the operation.
-Use cases for these operations are outlined in Table 1.
+deathtime=deadline+max_delay.deathtime=deadline+max\_delay. (1)
 The properties of the example <<Task>>s are as described:
+  • 1.
     effect – contains <<Task>>’s current position and state of <<Robot>> and <<Environment>> that the <<Task>> sets during execution or if finished.
 The examples are the <<Robot>>’s localisation or end-position of an object being transported,
+  • 2.
     deadline – the time during the <<Scenario>> by which we would like the <<Task>> to be completed,
+  • 3.
     priority – a priority of the <<Task>> (required by some <<DecAgent>>s), the higher, the more important the task is,
+  • 4.
     preemptive – defines if the <<Task>> can be interrupted after the <<Robot>> starts to work on it,
+  • 5.
     type – the name of the <<Task>> type in text form, useful for presenting the system’s state and scheduling algorithms assessments,
+  • 6.
     estimated_duration – estimated time required to complete the <<Task>> if its execution is started at the current time,
+  • 7.
     distance_from_robot – distance from the <<Robot>>’s position to the <<Task>>’s position,
+  • 8.
     deathtime – the time at which the task exceeds the deadline by maximum delay.
@@
 They must implement the following operations and communication interfaces.
-3.4.1 Configuring the Decision Agent
+2.3.1 Configuring the Decision Agents
 The <<DecAgent>> has one basic operation select_task() for selecting a <<Task>> for execution.
@@
 The functionality of the agents can be extended using <<AgentPlugin>>s.
+In one benchmark, you can execute different Decision Agents at once for comparison.
 2.3.2 Configuring the Evaluation Function
@@
 This is due to the infinite number of potential skills with which a <<DecAgent>> can be extended.
-<<DecAgent>>s wishing to use them must adapt to the way it works.
-By separating a capability as a <<AgentPlugin>>, there is no need to implement it individually for each <<DecAgent>>, and new <<DecAgent>>s are not forced to expand the code of the old ones to have it.
+By separating a capability as an <<AgentPlugin>>, there is no need to implement it individually for each <<DecAgent>>, and new <<DecAgent>>s are not forced to expand the code of the old ones to have it.
 Any <<DecAgent>> can use these <<AgentPlugin>>s.
@@
 For this reason, they require implementing a single update_job() operation (run within <<Scenario>>’s update_job() operation – Fig.
-4) that modifies the submitted <<Task>> based on the current time within the <<Scenario>>.
+3) that modifies the submitted <<Task>> based on the current time within the <<Scenario>>.
 By design, changes made by <<ScenarioPlugin>> should be characterised by significant complexity but potentially not desirable on every run.
@@
 We only require it to have an execute_step() operation that takes the currently selected <<Task>> to be executed and the <<Environment>> as arguments.
-3.5 Test-case configuration
+2.4 Test-case generators
+The test case configuration that is the occupancy map, human trajectories, and the real— unknown to the robot, task schedule, can be set manually for a given specific application, or can be generated with the prepared generators.
 2.4.1 The robot’s environment generation
@@
 The static obstacles emulating walls for the test session are generated using the recursive division method configured with the provided parameters (Fig.
-5).
+4(a)).
 First, all rooms, walls, and doors are generated considering the provided constraints.
@@
 Ultimately, we generate starting and destination poses for human actors and define their behaviour.
-Figure 5: Configurable environment parameters
+(a) Configurable environment parameters
+(b) Example randomly generated map with objects (blue), calculated probability distribution for pedestrian spawning (the darker green the bigger probability), and example footprints of the pedestrians (red) at the system runtime
+Figure 4: Robot’s environment configuration
 The probability map is calculated based on the pedestrian spawning probabilities set during the environment configuration (Fig.
-5/Pedestrians).
+4b/Pedestrians).
 Different probabilities can be set for the rooms, doors, and map entrance.
 The example environment map is presented in Fig.
-6.
+4(b).
 All of this satisfy [UC2.3], specifically [UC2.3.1] and [UC2.3.2].
-Figure 6: Sample randomly generated map with objects (blue), calculated probability distribution for pedestrian spawning (the darker green the bigger probability), and example footprints of the pedestrians (red) at the system runtime
 2.4.2 Task generation
@@
 For example, the <<Robot>>’s position should be able to navigate to, and the object’s position should be within one of the existing pieces of furniture.
-4 Example system
-4.1 System setup
+2.5 Example benchmarking sessions
+2.5.1 Decision agent training
+As mentioned before, DQN <<DecAgent>> requires a trained neural network to operate.
+Training this network may be considered an example of a successful benchmarking session utilising our Mobile Manipulator System.
+The <<DecAgent>> was taught using reinforcement learning (DQN algorithm).
+The training consisted of the <<DecAgent>> rehearsing numerous <<Scenario>>s.
+During the training, the <<Scenario>>s reflected 4 hours of social <<Robot>> work, during which the <<DecAgent>> was expected to complete 12 <<Task>>s of each of the 3 types.
+In each, the <<Task>>s were generated using a new seed.
+The DQN <<DecAgent>>’s actions were evaluated using DQN Eval, and the reward received during this stage was used as a reward for reinforcement learning.
+Learning took several million steps, each representing 5 seconds of virtual time.
+Individual <<Scenario>>s ended if all <<Task>>s were completed, one of the <<Task>>s died or <<Task>> completion exceeded the <<Scenario>> time.
+Once training was complete, the network was saved for further use.
+The training was repeated several times to verify differences in the performance of the different network structures and the value of the reward for individual actions.
+3 Illustrative example
 In the following section, we describe an example Mobile Manipulator System based on the <<TaBSASystem>>.
@@
 They are graphically presented in Fig.
-7.
-Figure 7: Mobile Manipulator System
-4.1.1 Tasks
-Mobile Manipulator Task is based on the <<Task>> stereotype but implements numerous additional operations and properties (see Fig. 8).
-Figure 8: Mobile Manipulator Task
-The Fall <<Task>> is meant to imitate the situation in which the <<Robot>> is informed about an elderly falling to the ground.
-The <<Robot>> moves to the chosen location and waits there for a given time duration.
-This imitates a brief examination of the elderly health.
-This <<Task>> has the highest ’priority’, cannot be interrupted, and is terminated about 15 minutes after its ’deadline’.
-The Patrol <<Task>> resembles the <<Robot>> moving down the path from the starting point to the end.
-This <<Task>> can be interrupted at any time and cannot die.
-Patrol’s ’priority’ is determined randomly when it is created, but it can never be higher than the Fall’s ’priority’ value.
-Pick <<Task>> and Place <<Task>> are not used on their own.
-They are responsible for the <<Robot>> picking up and placing the object.
-As our objects are represented merely by the points on the map, the <<Task>>s are performed by standing for a set amount of time near the object.
-These <<Task>>s received a ’priority’ of 0 for the purposes of the Pick And Place <<Task>> - they aren’t used independently, so they don’t need a realistic ’priority’ value, and we didn’t want their priority to influence complex <<Task>>s.
+5(a).
+(a) Mobile Manipulator System
+(b) Mobile Manipulator Task
+Figure 5: Mobile Manipulator system definition
+3.1 Tasks
+Mobile Manipulator Task is based on the <<Task>> stereotype but implements numerous additional operations and properties (see Fig. 5(b)).
 To represent tasks composed of several simpler tasks, we prepared the Complex <<Task>>.
@@
 The only composite <<Task>> prepared within the system was Pick And Place.
-It currently consists of 3 subtasks: Pick, Patrol, and Place.
 This <<Task>> aims to transport an item from one place to another.
-We set the Pick And Place <<Task>> as not interruptable.
-As these <<Task>>s utilize objects placed in the <<Environment>>, they tend to start and end close to each other.
-4.1.2 Decision agents
-We have implemented multiple <<DecAgent>>s to test the effectiveness of different <<Scenario>> execution approaches.
-Despite their inner workings, all <<DecAgent>>s respect the ’preemptive’ property of the <<Task>>s and won’t interrupt the <<Task>> that shouldn’t be paused during execution.
+3.2 Decision Agents
+  • 1.
     Simple-longest and Simple-shortest <<DecAgent>>s select the longest and the shortest <<Task>> for the execution, respectively.
-Additionally, they possess the ’hesitance’ property.
-Every time the <<Task>> is selected for execution, a random number between 0 and 1 is generated.
-If it is smaller than the hesitance value, the <<DecAgent>> returns the same <<Task>> as in the previous decision without checking anything (if the <<Task>> is still within the received ’jobs’).
+  • 2.
     The Distance <<DecAgent>> chooses the <<Task>> that is the furthest from the <<Robot>>.
-Additionally, the <<DecAgent>> possesses the ratio property used to offset the distance by a fraction of the <<Task>>’s estimated duration.
-If the ’ratio’ is non-zero, the <<DecAgent>> effectively chooses the <<Task>> with the highest ’dist_score’ calculated by subtracting the ’ratio’ times the <<Task>>’s ’estimated_duration’ from the <<Task>>’s ’distance_from_robot’:
-$dist\_score=$ $\max_{jobs}(job.distance\_from\_robot-ratio*$ (2)
-$*job.estimated\_duration)$
-That means that a higher ratio causes the <<DecAgent>> to favour shorter <<Task>>s.
+  • 3.
     The Scheduler <<DecAgent>> uses the Request Table <<AgentPlugin>>.
@@
 <<Task>>s with the same ’priority’ are then sorted by the earliest ’request_time’.
-Then the <<AgentPlugin>> creates a schedule of sorts using <<Task>>s’ ’deadlines’ and ’estimated_durations’, returning the lists of scheduled and rejected <<Task>>s.
-Going through the sorted list, if the <<Task>>’s time from start to finish collides with any other <<Task>> in the scheduled list, the <<Task>> is assigned to the rejected list and vice versa.
-In the end, the Scheduler <<DecAgent>> receives both of the lists.
-It returns the <<Task>> from the scheduled list, which time of execution from start to finish contains the current time.
-If there is no such <<Task>>, the <<DecAgent>> doesn’t return any of them.
-The <<Task>>s which aren’t worked on at their scheduled time will not be completed in the free time.
-The last <<DecAgent>> is called the DQN because it uses the neural network trained using the Deep Q Network [26] (DQN) algorithm to choose the <<Task>> for execution.
-The network’s input is of the shape [number of <<Task>> types][number of <<Task>>s per type][3] as presented in Fig.
-9.
-The proposed system also contains the interface for the RLlib library [22].
-The RLlib contains multiple Reinforcement Learning (RL) state-of-the-art methods that can also be tested for training and benchmarking algorithms scheduling tasks for mobile robots working in dynamic environments, so alternatively, many other RL algorithms can be used in this block.
-Considering that only Fall, Patrol, and Pick And Place <<Task>>s are directly used within the Service Scenario, in our case, the number of <<Task>> types is 3.
-For each <<Task>> type, the set number of <<Task>>s is selected by the earliest deadline.
-Sometimes, the amount of <<Task>> may be too low to fill the whole input.
-In such cases, the empty places are filled with zeros.
-For each <<Task>>, three properties are placed inside the network’s input: ’estimated_duration’, ’preemptive’ (represented as integer), and ’distance_from_robot’.
-<<Task>>s placed within the network’s input are remembered as a list.
-The network processes the input and outputs the values for each <<Task>>.
-The highest value marks the position of the <<Task>> to be executed within the saved list.
-The important process of training the DQN <<DecAgent>> (as per [UC2]) takes place inside its select_task() operation using deep reinforcement learning.
-After the <<DecAgent>> calculates the new input (shape presented in Fig.
-9) it can supply it along with the previous input, selected <<Task>>, and ’reward’ from the <<EvalFunction>> to the training memory used to teach the network.
-Training the <<DecAgent>> is a long process and requires running multiple <<Scenario>>s using the same <<DecAgent>> (as presented in Fig.
-3) as per [UC2.4].
-Figure 9: DQNAgent’s network input and output
-Apart from the Request Table, our system also contains another <<AgentPlugin>> – Task Predictor.
-Its purpose is to predict <<Task>>s that can be requested in the near future before they are called.
-This <<AgentPlugin>> possesses an instance of the Request Table <<AgentPlugin>>.
-Its most important part is a neural network with Long Short-Term Memory (LSTM) network [31] layers.
-The plugin is meant to predict <<Task>>s within the set time horizon divided into slots.
-For each slot, 3 values are calculated: number of <<Task>>s within the slot, number of <<Task>>s set to be executed within the slot (per information from Request Table), and the sum of the ’priorities’ of executable <<Task>>s within the slot.
-These values, along with the seed used for <<Task>> generation, are passed to the network.
-The network output consists of the same three values for the same amount of slots but for time horizon counted since one slot in the future.
-The visualisation of this network’s input and output can be found in Fig.
-10.
-Calculating the differences between the output and the input allows predicting if new <<Task>>s will be added within these time slots in the future.
-Figure 10: Task Predictor’s network input and output
-4.1.3 Scenario configurations
-The example Service Scenario, based on the <<Scenario>> stereotype, has a Mobile Manipulator <<Robot>>, a single <<ScenarioPlugin>> – Duration Estimator Network, and only Mobile Manipulator Task <<Task>>s (Fig.
-11).
+  • 4.
+    DQN <<DecAgent>> implements the neural network trained using the Deep Q Network [9] (DQN) algorithm to choose the <<Task>> for execution.
+3.3 Scenario
+The example Service Scenario, based on the <<Scenario>> stereotype, has a Mobile Manipulator <<Robot>>, a single <<ScenarioPlugin>> – Duration Estimator Network, and only Mobile Manipulator Task <<Task>>s.
 The <<Scenario>> is meant to represent several hours of work for the mobile service <<Robot>>.
@@
 The properties of these <<Task>>s are updated every time by the update_job() operation: estimate_duration() is called, ’distance_from_robot’ is calculated, and every path that needs to be planned within the <<Task>> is refreshed using the current state of the environment.
-Figure 11: Service Scenario
-Duration Estimator Network <<ScenarioPlugin>> is as the name suggests meant to estimate the duration of the <<Task>>.
-Despite every <<Task>> in our Mobile Manipulator System having its own estimate, in the real world, these numbers may be offset by an inconsistent environment caused by, for example, humans walking around or misplacement of furniture and objects.
-Therefore precise prediction of the <<Task>>’s duration is impossible.
-The plugin consists of a neural network trained offline on data gathered from multiple runs.
-To mitigate overtraining the network, during data gathering, ’durations’ and ’deadlines’ of <<Task>>s are slightly modified by adding noise and additional <<Task>>s are added randomly to the <<Scenario>>.
-The network has a single output representing the predicted <<Task>> duration.
-Its input, on the other hand, consists of the following values: the day when the predicted <<Task>> is performed (seed from <<Task>> configuration), current time, <<Task>>’s ’deadline’, <<Task>>’s current position, <<Task>>’s goal position (if it requires movement, otherwise current position again), <<Task>>’s current ’distance_from_robot’ and <<Task>>’s ’priority’.
-The structure is presented in Fig.
-12.
-Figure 12: Duration Estimator Network’s network input and output
-4.1.4 Evaluation functions
+3.4 Evaluation functions
 Two <<EvalFunction>>s are currently implemented within the Mobile Manipulator System: DQN Eval and Statistic Eval. Both of them have their specific purposes.
@@
 The additional parameter— ’penalty_change_job’ is added to the reward only in the last two cases if the current action is different than the ’previous_action’ property.
-$\footnotesize reward=\left\{\begin{array}[]{ll}reward\_all\_complete&if\,all\,tasks\,are\\ &\quad completed\\ penalty\_dead\_job&if\,at\,least\,one\,job\\ &\quad is\,dead\\ reward\_job\_complete&if\,job\,was\,just\\ &\quad completed\\ 0&if\,there\,is\,no\,jobs\\ &\quad in\,the\,list\\ reward\_real\_job&if\,real\,job\,was\\ &\quad selected\\ penalty\_nonexistent\_job&if\,nonexistent\,job\\ &\quad was\,chosen\\ \end{array}\right.$ (3)
+[…]
 Statistic Eval aims to numerically assess the work of any <<DecAgent>> it monitors using statistics.
@@
 Statistic Eval returns the ’terminate’ flag if all <<Task>>s are completed, one of the <<Task>>s is terminated or the <<DecAgent>> switches between <<Task>>s and returns to the same one for the third time within 3 minutes.
-Other useful values returned by the evaluation function on each iteration are represented by the following properties:
-  • full_travel_distance – the full travel distance of the <<Robot>> calculated by comparing the <<Robot>>’s current pose to the remembered ’last_robot_pos’,
-  • num_of_tasks_completed – the number of completed <<Task>>s of every type in ’task_types’,
-  • task_completion_to_deadline – the difference between the <<Task>>’s completion time and the original ’deadline’,
-  • task_completion_to_deathtime – the difference between the <<Task>>’s completion time and potential time of termination, None if a <<Task>> can’t be terminated,
-  • num_of_tasks_interrupted – number of <<Task>>s’ interruptions per <<Task>> type in ’task_types’,
-  • task_interruptions – number of <<Task>>s’ interruptions for each <<Task>> in <<Scenario>>’s ’task’ list,
-  • num_of_human_abandonement – number of <<Task>>s’ interruptions for each <<Task>> and number of times the <<Robot>> has abandoned a human in need.
-The last statistic is specific to our system and represents the number of situations where the <<Robot>> drove within ’abandonment_distance’ of any Fall <<Task>> without working on it.
-4.2 Example benchmarking sessions
-4.2.1 Decision agent training
-As mentioned before, DQN <<DecAgent>> requires a trained neural network to operate.
-Training this network may be considered an example of a successful benchmarking session utilising our Mobile Manipulator System.
-The <<DecAgent>> was taught using reinforcement learning (DQN algorithm).
-The training consisted of the <<DecAgent>> rehearsing numerous <<Scenario>>s.
-During the training, the <<Scenario>>s reflected 4 hours of social <<Robot>> work, during which the <<DecAgent>> was expected to complete 12 <<Task>>s of each of the 3 types.
-In each, the <<Task>>s were generated using a new seed.
-The DQN <<DecAgent>>’s actions were evaluated using DQN Eval, and the reward received during this stage was used as a reward for reinforcement learning.
-Learning took several million steps, each representing 5 seconds of virtual time.
-Individual <<Scenario>>s ended if all <<Task>>s were completed, one of the <<Task>>s died or <<Task>> completion exceeded the <<Scenario>> time.
-Once training was complete, the network was saved for further use.
-During the project, the training was repeated several times to verify differences in the performance of the different network structures and the value of the reward for individual actions.
-4.2.2 Agent comparison
-Another example of a benchmarking session could be an attempt to compare the performance of individual decision-making <<DecAgent>>s.
-Such an exercise makes it possible to determine the best one and to identify potential errors in its performance.
+In the documentation, there are other useful values returned by the evaluation function on each iteration.
+3.5 Experimental results
 Initially, 50 <<Scenario>>s with random <<Task>>s were generated for each <<DecAgent>> under test.
@@
 Scheduler and Simple-longest <<DecAgent>>s both failed for different reasons.
-If two Fall <<Task>>s were to be scheduled at the same time, Scheduler would only complete one of them.
-Simple-longest agent, on the other hand, prioritizes long <<Task>>s, which causes him to start with Pick And Place <<Task>>s, which tend to be longer than the others.
-DQN <<DecAgent>> didn’t have a good outcome during this test either, but something else was also spotted.
-After training the first DQN <<DecAgent>>, a lot of <<Scenario>>s ended due to running out of time.
+During the framework execution, we identified unwanted behaviour of the first DQN <<DecAgent>>.
+It has terminated a lot of <<Scenario>>s due to running out of time.
 This fact prompted us to reevaluate the code and discover that the penalty for doing nothing was indeed wrongly applied as a positive number.
@@
 Fig.
-14 shows the absolute difference between <<Task>>s’ completion times and their original ’deadlines’ sorted by <<Task>> type.
+7(a) shows the absolute difference between <<Task>>s’ completion times and their original ’deadlines’ sorted by <<Task>> type.
 As can be seen, one of the <<DecAgent>>s seems to be much better than the others - Scheduler - as it displayed much smaller disparities.
 This means that <<Task>>s worked on by this agent were completed close to the original ’deadlines’, which may be a desirable behaviour in some cases.
-All of the other <<DecAgent>>s start performing <<Task>>s as fast as they come, which causes the differences to be quite large.
-We can also observe that the biggest differences are visible for Pick And Place <<Task>>s, which suggests that they are usually the longest to complete or are susceptible to moving pedestrians.
-Figure 14: Difference between real completion time and the ’deadline’ for different <<Task>> types for different <<DecAgent>>s
+(a) Difference between real completion time and the ’deadline’ for different <<Task>> types for different <<DecAgent>>s
+(b) Number of completed <<Task>>s per type for different <<DecAgent>>s
+Figure 7: Comparison of the scheduling algorithms
 Fig.
-15 shows the number of <<Task>>s of each type completed by each <<DecAgent>> during the mentioned 50 <<Scenario>>.
+7(b) shows the number of <<Task>>s of each type completed by each <<DecAgent>> during the 50 <<Scenario>>s.
 These plots also confirm that most <<DecAgent>>s work as intended despite failing the <<DecAgent>>s.
@@
 The biggest surprise is the first DQN, which seems to have completed almost as many fall <<Task>>s as the Distance <<DecAgent>> but didn’t complete any <<Scenario>>s.
-This further confirms the need to retrain the <<DecAgent>> with different parameters.
-Figure 15: Number of completed <<Task>>s per type for different <<DecAgent>>s
+While the DQN-based decision agent performed poorly in our experiments, this outcome highlights a central feature of the proposed framework: its ability to reveal limitations of specific approaches under controlled, reproducible conditions.
+The framework is designed precisely for such evaluations, enabling algorithm developers to identify, analyze, and address weaknesses in their methods.
+Thus, the weaker performance of DQN in this setting demonstrates the effectiveness of the proposed framework in revealing algorithmic limitations or providing valuable insights for further improvements.
 After that, another experiment was performed, in which each <<DecAgent>> completed the same 100 <<Scenario>>s (same seed used for <<Task>> generation).
 This allowed us to inspect the differences in <<DecAgent>>s’ behaviours closely.
-Figures 16 and 17 show how the same scenario was completed by Simple-shortest and Distance <<DecAgent>>s.
+Figures 8(a) and 8(b) show how the same scenario was completed by Simple-shortest and Distance <<DecAgent>>s.
 The horizontal axis represents the passage of time, while the currently performed <<Task>> is identified by its colour and order in the <<Task>> list of the <<Scenario>>.
@@
 On the other hand, since about 1700th second, the outcomes are almost identical, probably because these <<Task>>s were given to the agents at the same time late into the day.
-Figure 16: Scenario course as performed by Simple-shortest <<DecAgent>>
-Figure 17: Scenario course as performed by Distance <<DecAgent>>
-5 Summary
-This article tackles the problem of assessing task-scheduling algorithms in the mobile robot domain.
-We propose a benchmarking framework that encapsulates the scheduling algorithm with the <<DecAgent>> interface.
-Numerous algorithms are optimised for given robot applications, robot types, or environments.
-Therefore, our framework allows the configuration of the robot’s kinematics, tasks, environment, and testing scenarios.
-By the configuration, the framework reflects the desired robot system and its environment setup.
-The system is open-source, and its repository is shared [^2] https://github.com/RCPRG-ros-pkg/Smit-Sim.
-Therefore, our experiments can be easily re-executed, and the framework can be used to evaluate new or well-known algorithms in the domain of mobile robots.
-The framework incorporates the generators for schedules of <<Task>> requests, trajectories of moving humans, and maps with objects for manipulation and rooms.
-These generators can be used in training task-scheduling algorithms and for statistic-based evaluation of these or other algorithms.
-We have validated the framework by configuring it for a mobile service <<Robot>> application, teaching AI-based algorithms using DQN, and assessing it against basic scheduling algorithms.
-As a result of the framework application, we could identify flaws in both coding and decision-making.
-Ultimately, we could understand and graphically present the core differences between the selected algorithms at chosen configurations.
-In the given test case, scheduling algorithms that chose the shortest (Simple-shortest <<DecAgent>> with ’hesitance’ 0.5) or the furthest (Distance <<DecAgent>> with ’ratio’ 0.5) task managed to complete the <<Scenario>>s respectively  90% and  60% of the time.
-They also didn’t seem to favour any specific type of <<Task>>.
-The algorithm choosing the longest <<Task>> (Simple-longest<<DecAgent>> with ’hesitance’ 0.5) did the worst (with all <<Scenario>>s fatally failed), followed closely by the AI-based ones (both versions of DQN <<DecAgent>>) and the one based on the strict schedule (Scheduler <<DecAgent>>).
-In the first case, the <<DecAgent>> favours executing Pick And Place, while the others Fall <<Task>>s.
-Of all the algorithms, only the schedule-based one completed the <<Task>>s almost at the set ’deadline’.
-From testing the algorithms in the proposed benchmarking system, we not only select the best algorithm but also learn about the <<Robot>> application.
-During the investigation of our example, we learned that both of the best algorithms managed to get ahead by completing as many <<Task>>s as fast as possible by favouring shorter <<Task>>s.
-Additionally, the Distance <<DecAgent>> avoided tasks that are close to each other.
-Pick And Place <<Task>>s tend to start and end close to each other and are often the longest to perform, so algorithms that favour them for any reason are likely to fail.
-Seeing that most scenarios ended due to the "dead" status (which happens when Fall <<Task>> is not completed for some time after its ’deadline’), we find that keeping deadlines is also critical in this application.
-Thus, the scheduling algorithms that are aware of deadlines should perform better than the ones that are not.
-On the other hand, a <<DecAgent>> may concentrate on Fall <<Task>>s but can still fail if it uses a subpar scheduling algorithm.
-Scheduling tasks of mobile robots is challenging in a dynamic environment.
-However, based on the experience gained in this research, we find the advanced scheduling of manipulation tasks the next valuable target.
-Advanced scheduling must involve semantic planning of the tasks, which is a state-of-the-art solution to managing complexity in robot tasks [29].
-[…]
+(a) Scenario course as performed by Simple-shortest <<DecAgent>>
+(b) Scenario course as performed by Distance <<DecAgent>>
+Figure 8: Comparison of the example scenario execution
+4 Impact
+The presented framework has demonstrated significant utility in the quantitative evaluation of scheduling algorithms, referred to as Decision Agents, for service robots operating in dynamic environments.
+By executing each agent within identical, simulated scenarios that include randomized human movement and diverse tasks, the framework enables fair, reproducible comparisons.
+It supports both broad benchmarking across general conditions and targeted testing in application-specific setups.
+The experiments confirmed the framework’s capability to reveal subtle behavioral differences between agents, highlight performance patterns, and, crucially, expose implementation or configuration errors—such as misapplied penalties in one of the tested DQN agents.
+Through detailed statistical outputs and visual comparisons, users can assess algorithm versatility or select the most suitable scheduling strategy for a given robotic application, thereby accelerating both algorithm development and deployment-readiness.
+5 Conclusions
+We proposed and validated a practical benchmarking framework for evaluating scheduling algorithms in mobile service robotic applications.
+The framework enables configurable simulation of environments, tasks, robot kinematics, and human activity, allowing reproducible and comparative testing of decision-making algorithms under identical conditions.
+The experiments demonstrated the framework’s effectiveness in identifying both behavioral characteristics and implementation flaws, highlighting its diagnostic value.
+By visualizing differences in task execution, completion timing, and failure causes, the framework supports both algorithm tuning and application-driven selection.
+The presented results confirm that the framework can effectively benchmark diverse decision-making agents, from heuristic schedulers to learning-based approaches.
+The fact that certain algorithms, such as DQN, underperformed demonstrates the framework’s usefulness.
+It provides a fair and transparent way to assess algorithms and to identify cases where they can be further refined.
+It is open-source and extensible [^2] https://github.com/RCPRG-ros-pkg/Smit-Sim, therefore it serves as a ready-to-use tool for practitioners aiming to improve scheduling strategies or adapt them to specific robot deployments.
+Finally, the insights gained suggest that task-deadline awareness and preference for shorter tasks increase scheduling robustness in realistic scenarios, while future work should target semantic-level planning for more complex robot activities.
 Acknowledgements